OpenStudy (anonymous):

please solve it \[\int\limits_{0}^{\pi}\sin (x) \cos(nx) dx\]

OpenStudy (anonymous):

answer is \[-2\div (n^{2}-1) \] ,for n=even and 0 for n= odd

OpenStudy (zarkon):

use integration by parts twice

OpenStudy (unklerhaukus):

alternatively use this trig identity\[\sin (u) \cos (v)=\frac 12\left[ \sin(u+v)+\sin(u-v)\right]\] \[\sin (x) \cos (nx)=\frac 12\left[ \sin((n+1)x)+\sin((1-n)x)\right]\]

OpenStudy (anonymous):

good